# Adaptive Text Watermark for Large Language Models

### v1

@@ 3.2 Semantic-based Logits Scaling Vector Extraction
The Semantic-based Logits Scaling Vector Extraction (see, Figure 1) is primarily developed to ensure the security of the watermark. The intuition of this approach is to make sure that the perturbation in the distribution is determined by the semantics of the generated text using neural networks so that the inserted watermarking cannot be easily decrypted by attackers.
@@
0:  Sentence Embedding Model (SE), Semantic Mapping Model (SM), Input sentence $S$.
@@
In particular, for tokens with distributions satisfying $H(P_{MM}(\cdot|S_{0:t-1}))\geq\alpha$ at time step $t$, we proceed to extract logits scaling vector based on the semantics of generated text. First, we obtain the sentence embedding using a Sentence Embedding Model (SE), i.e., $u^{(t)}(S)={\rm SE}({S}_{0:t-1})$, where $u^{(t)}\in\mathbbm{R}^{L}$. We employ a pre-trained LM as the Sentence Embedding Model, e.g., Sentence-BERT (Reimers & Gurevych, 2019), to extract the sentence embedding. Then, we train a Semantic Mapping Model (SM) to convert the sentence embedding into a logits scaling vector, denoted as $v^{(t)}(S)={\rm SM}(u^{(t)}(S);\theta)$, where $v^{(t)}\in\mathbbm{R}^{|\mathcal{V}|}$ and $\theta$ denotes the parameters of SM.
Specifically, we train a feedforward neural network with multiple residual connections so that the logits scaling vector should satisfy the following properties.
(1) Smoothness with respect to semantics: As we need to recover the logits scaling vector to perform detection even with modifications in the watermarked text, SM needs to be insensitive to small changes in semantics to improve the robustness of the watermark. In order to make the logits scaling vector aligns with the change in semantic embedding, we want to make sure that if two different sentences $S$ and $S^{\prime}$ are close in embedding space, they will have similar logits scaling vector $v$ and $v^{\prime}$. We will minimize the loss $\big{|}D(u(S),u(S^{\prime}))-D(v(S),v(S^{\prime}))\big{|}$, where we use the Euclidean distance $D(a,b)=\sqrt{\sum_{i=1}(a_{i}-b_{i})^{2}}$ to measure the distance between two sentence embeddings.
(2) Uniform Perturbation: To make the watermark easy to detect, we want to ensure the logits scaling vector increases the probability of roughly half of the tokens in the vocabulary. For any sentence $S$ within the training dataset $\mathcal{T}$, we will ensure that $\sum_{i\in\mathcal{V}}\mathrm{sign}\big{(}v_{i}(S)\big{)}=0$. Therefore, an equal rate of positive and negative entries.
@@
Therefore, considering all the above properties, the loss function of the Semantic Mapping Model is:
$\mathcal{L}(\mathcal{T},\theta)=$ $\sum_{S,S^{\prime}\in\mathcal{T}}\big{|}D(u(S),u(S^{\prime}))-D(v(S),v(S^{\prime}))\big{|}$ (2)
$+$ $\sum_{S\in\mathcal{T}}\big{|}\sum_{i\in|\mathcal{V}|}\mathrm{sign}(v_{i}(S))\big{|}+\sum_{i\in|\mathcal{V}|}\big{|}\sum_{S\in\mathcal{T}}\mathrm{sign}(v_{i}(S))\big{|}.$
@@
3.3 Adaptive Watermark Temperature Scaling
To further minimize the impact of perturbation on the original probability distribution, we introduce the Adaptive Watermark Temperature Scaling method (see, Figure 1). This technique enhances the consistency of the perturbation with the original distribution by adaptively adjusting the temperature of the distribution.
@@ 3.3 Adaptive Watermark Temperature Scaling
where $\delta>0$ controls the strength of the watermark. In this way, we adaptively perturb the logit of each token by proportionally scaling up a factor of $(1+\delta\cdot\hat{v}^{(t)})$ using the obtained SLSVE. If the original logit $l_{i}^{(t)}$ is small, the $\hat{l}_{i}^{(t)}$ after the manipulation remains small, indicating that our watermark has a minimal effect on these logits. Our temperature scaling mainly affects tokens with large original logits $l_{i}^{(t)}$, causing the corresponding $\hat{l}_{i}^{(t)}$ to increase more significantly. In contrast to uniformly increasing the logits of tokens in the ‘green list’ as in previous work (Kirchenbauer et al., 2023a), our approach is to embed a watermark by modulating the temperature of the logits. Therefore, our watermark strength relies on the original distribution, which will make the generated watermarked text more coherent. As we will show in Figure 4, AWTS also contributes to the improvement of text quality.
@@
Algorithm 2 describes the whole process of our Adaptive Watermark algorithm. At the beginning of the text generation process, only a limited number of tokens can be used for extracting semantic embeddings, which makes the detection of the watermark unstable under attack. To maintain the stability and robustness of the watermark, we set a secret opening sentence $\hat{S}$. When the number of generated tokens is smaller than a pre-determined measure threshold $M$, we extract the semantic embedding from the opening sentence and apply SLSVE.
@@ 4 Adaptive Watermark Detection
In this paper, we propose to approximate the Likelihood Ratio Test (LRT) to detect watermark. Consider the following null hypothesis $H_{0}$ and alternative hypothesis $H_{1}$,
@@
Given a sequence of tokens $S=\{S^{(1)},S^{(2)},...,S^{(T)}\}$, we input the sequence to MM to calculate the entropy and select the potential watermarked token at each time step, following the same process as the AWTI described in Section 3.1. Denote the set of potential watermarked token as $\mathcal{W}$, i.e., for all $t\in\mathcal{W}$, $H(P_{MM}(\cdot|S_{0:t-1}))\geq\alpha$ holds. For any $t\in\mathcal{W}$, if it is sampled from the perturbed distribution, the distribution would be $P_{H_{1}}(\cdot|R,S_{0:t})$. Conversely, if it is sampled from the original distribution, the distribution would be $P_{H_{0}}(\cdot|R,S_{0:t})$. Subsequently, the logits scaling vector is extracted utilizing both the Semantic Embedding Model and the Semantic Mapping Model, as detailed in Section 3.2. As we perturb the distribution with temperature scaling, the likelihood ratio can be written as
@@
where we omitted the superscript of time $t$ and the normalization factor in the softmax function.
Figure 2: Workflow of Adaptive Detection. This figure illustrates a single time step during the detection process. At each time step, we will first check if the current token is the potential watermarked token by applying AWTI to assess the entropy of the current distribution. For potential watermarked token, the preceding text will be used to extract the logits scaling vector through SLSVE. If the corresponding value of the potential watermarked token in the logits scaling vector is positive, it will be added to the total score.
@@
Score(${\rm Score}(S=\{s^{(1)},...,s^{(T)}\})=\frac{\sum_{t\in W}\delta\cdot\hat{v}^{(t)}_{s^{(t)}}}{|W|},$ (5)
where $|W|$ represents the number of all potential watermarked tokens.
5 Experiment
@@ 5.1 Experiment Setting
Dataset and Prompt. The realnewslike subset in C4 dataset (Raffel et al., 2020) is used to test the performance of our methods. We use the first two sentences in each text as a prompt for LLMs, and the subsequent $200$ tokens as the human-generated text. We also generate $200\pm 30$ tokens using the LLMs and prompts. The Semantic Mapping Model is trained using the Sentence-Compression dataset  [^1] https://huggingface.co/datasets/embedding-data/sentence-compression.
Baseline and Language Model. We apply our method to three different models, which are OPT-2.7B, OPT-6.7B (Zhang et al., 2022) and GPT-J-6B (Wang & Komatsuzaki, 2021). At each time step, the next token is sampled from the distribution using a combination of Top-$K$ and Top-$p$ sampling methods, with $K=50$ and $p=0.9$. In our watermarking algorithm, the default hyperparameters are set as follows: $\alpha=2$, $\delta=1.5$, $M=50$. We use Sentence-Transformers (Reimers & Gurevych, 2019) to extract the sentence embedding. The specific model we used is all-mpnet-base-v2  [^2] https://huggingface.co/sentence-transformers/all-mpnet-base-v2. Our method is compared with existing methods, including KGW-00 (Zhao et al., 2023), KGW-$1$/KGW-$2$ (Kirchenbauer et al., 2023a), EXP-edit (Kuditipudi et al., 2023). In KWG-$1$/KWG-$2$, the numeral denotes the hashing of the previous 1 or 2 tokens, respectively. For KGW-00, KGW-$1$, and KGW-$2$, the ‘green list’ token percentage is set to be $0.5$.
Evaluation Methods. Watermark detection performance is evaluated through ROC-AUC value and the Best F1 Score. In addition, We present the ROC curve, which illustrates the true positive rate (TPR) and false positive rate (FPR) at various thresholds. The TPR and FPR indicate the accurate identification of watermarked text and the incorrect recognition of human-written text as being watermarked. The perplexity of the text is calculated using LLama-13B (Touvron et al., 2023), measuring the quality of the text.
@@ 5.2 Robustness
To demonstrate the robustness of our watermark strategy, we execute a paraphrase attack on the watermarked text using two different paraphrasing methods: GPT-$3.5$ and DIPPER (Krishna et al., 2023). For GPT-$3.5$, we use the gpt-3.5-turbo-instruct version, and we employ the same prompt as used in Kirchenbauer et al. (2023b) to paraphrase the watermarked text. DIPPER is a paraphrase generation model to rewrite the text and evade the watermark detection, and the lexical diversity is set to be $60$.
The results are summarized in Table 1, which shows that all methods demonstrate effective performance when the watermarked texts remain unaltered. However, when the watermarked text is paraphrased using GPT-$3.5$ and DIPPER, for OPT-$6.7$, the ROC-AUCs of our method are $0.981$ and $0.964$, respectively. Additionally, the best F1 score achieves $0.932$ and $0.903$. This performance significantly exceeds all the compared baselines. The experimental results on OPT-2.7 and GPT-J-6B also show that our watermark method is more robust than other methods under paraphrase attack.
@@ 5.3 Text Quality
The quality of the generated text is evaluated by perplexity. A lower perplexity value suggests better text quality. As shown in Figure 3, we compare the perplexity of text generated by our method with human-written text, un-watermarked text (texts generated by the model without any watermarking), and other baseline methods. Our method has a negligible impact on the perplexity compared to the un-watermarked text. It also shows comparable results with EXP-edit, which is a distortion-free watermark method. The quality of text watermarked using our approach is clearly superior to that of KGW-00 and KGW-$1$.
@@ 5.4 Security Results
The spoofing attack (Sadasivan et al., 2023) is used to evaluate the security of our method, which is an adversarial attack designed to produce non-AI text that can be mistakenly identified as AI-generated by the watermark detector. This is accomplished by frequently querying the watermarked language model and examining the frequency of the 181 most commonly used words within the vocabulary. For our method, we use the sentences in the DBPedia Classes Dataset  [^3] https://www.kaggle.com/datasets/danofer/dbpedia-classes/data as prompt to generate watermarked text and analyze the word frequency. DBPedia Classes Dataset has three levels of hierarchy (L1, L2, L3). From L1 to L3, the granularity of classification becomes increasingly detailed, and the range of contexts within a specific class narrows down for different sentences. We compare the decryption rate among different watermark methods, which is the percentage of ‘green’ list tokens that the attacker can decrypt. From Table 2, we can observe that KGW-0 exhibits the highest decryption rate because of its fixed ’green/red’ list. For KGW-1 and KGW-2, the implementation of the previous token as a hash key adds a certain level of difficulty to the decryption process compared to the KGW-0. Our method demonstrates superior resistance to attack compared to those works, with a decryption rate of only $0.571$.
@@
Language Model Evaluation KGW-0 KGW-1 KGW-2 Ours (L3)

### v2

@@ 3.2 Semantic-based Logits Scaling Vector Extraction
The Semantic-based Logits Scaling Vector Extraction (see Figure 1-➁) is primarily developed to ensure the security of the watermark. The intuition of this approach is to make sure that the perturbation in the distribution is determined by the semantics of the generated text using a well-designed neural network so that the inserted watermarking cannot be easily decrypted by attackers.
@@
0:  Sentence Embedding Model ($\rm SE$), Semantic Mapping Model ($\rm SM$), Input sentence $S$.
@@
In particular, for tokens with distributions satisfying $H(P_{MM}(\cdot|S_{0:t-1}))\geq\alpha$ at step $t$, we proceed to extract logits scaling vector based on the semantics of generated text. First, we obtain the semantic embedding using a Sentence Embedding ($\rm SE$) Model, i.e., $u^{(t)}(S)={\rm SE}({S}_{0:t-1})$, where $u^{(t)}\in\mathbbm{R}^{L}$. We employ a pre-trained LM as the Sentence Embedding Model, e.g., Sentence-BERT (Reimers & Gurevych, 2019), to extract the sentence embedding. Then, we train a Semantic Mapping ($\rm SM$) Model to convert the sentence embedding into a logits scaling vector, denoted as $v^{(t)}(S)={\rm SM}(u^{(t)}(S);\theta)$, where $v^{(t)}\in\mathbbm{R}^{|\mathcal{V}|}$ and $\theta$ denotes the parameters of the Semantic Mapping Model.
Specifically, we train a feedforward neural network with multiple residual connections to ensure that the logits scaling vector satisfies the following properties.
(1) Smoothness with respect to semantics: As we need to recover the logits scaling vector to perform detection even with modifications in the watermarked text, the Semantic Mapping Model needs to be insensitive to small changes in semantics to improve the robustness of the watermark. In order to make the logits scaling vector align with the change in semantic embedding, we want to make sure that if two different sentences $S$ and $S^{\prime}$ are close in embedding space, they will have similar logits scaling vector $v$ and $v^{\prime}$. We will minimize the loss $\big{|}D(u(S),u(S^{\prime}))-D(v(S),v(S^{\prime}))\big{|}$, where we use the Euclidean distance $D(a,b)=\sqrt{\sum_{i=1}(a_{i}-b_{i})^{2}}$ to measure the distance between two sentence embeddings.
(2) Uniform Perturbation: To make the watermark easy to detect, we want to ensure the logits scaling vector increases the probability of roughly half of the tokens in the vocabulary. For any sentence $S$ within the training dataset $\mathcal{T}$, we ensure that $\sum_{i\in\mathcal{V}}\mathrm{sign}\big{(}v_{i}(S)\big{)}=0$. Therefore, an equal rate of positive and negative entries in $v_{i}(S)$.
@@
Additionally, during the training of the Semantic Mapping Model, we ensure that the logits scaling vector remains stable against semantic variations using the following three implementation tricks. This further guarantees that the watermark’s robustness remains unaffected by semantic variations within a specific range. (i) Due to the high-dimensional nature of sentence embeddings, experiments on the original training dataset revealed that most sentence embeddings are essentially orthogonal to each other. To improve the model’s capability in effectively handling slight semantic perturbations, we enlarged our training dataset by performing data augmentation, including paraphrasing, expanding, and shortening the sentences in the original training dataset. (ii) With the augmented data mentioned above, a contrastive loss term is applied during training to minimize the distance between the sentence ($S\in\mathcal{T}$) and its augmented counterparts($\widetilde{S}\in\widetilde{\mathcal{T}}$), i.e., $\min\sum_{S\in\mathcal{T},\widetilde{S}\in\widetilde{\mathcal{T}}}\big{|}D(v(S),v(\widetilde{S}))\big{|}$. This contrastive term improves the stability of the model when a modified sentence is presented. (iii) We rescale the Euclidean distance between two sentence embeddings $d=D(u(S),u(S^{\prime}))\in[L,U]$ to a wider range $d^{\prime}\in[L^{\prime},U^{\prime}]$ using a linear transformation, where $L^{\prime}<L$ and $U^{\prime}>U$. Rescaling makes the distance between two sentence embeddings smaller when they are close to each other and increases the distance when they are originally farther apart in the embedding space.
Therefore, considering all the desired properties, the loss function of the Semantic Mapping Model is:
$\mathcal{L}(\widetilde{\mathcal{T}},\theta)=$ $\sum_{S,S^{\prime}\in\widetilde{\mathcal{T}}}\!\big{|}D(u(S),u(S^{\prime}))\!-\!D(v(S),v(S^{\prime}))\big{|}$
$+$ $\!\sum_{S\in\widetilde{\mathcal{T}}}\!\big{|}\!\sum_{i\in|\mathcal{V}|}\mathrm{sign}(v_{i}(S))\big{|}\!+\!\sum_{i\in|\mathcal{V}|}\!\big{|}\!\sum_{S\in\widetilde{\mathcal{T}}}\mathrm{sign}(v_{i}(S))\big{|}$
$+$ $\sum_{S\in\mathcal{T},\widetilde{S}\in\widetilde{\mathcal{T}}}\big{|}D(v(S),v(\widetilde{S}))\big{|}.$ (2)
@@
3.3 Adaptive Watermark Temperature Scaling (AWTS)
To further minimize the impact of perturbation on the original probability distribution, we introduce the Adaptive Watermark Temperature Scaling method (see Figure 1-➂). This technique enhances the consistency of the perturbed distribution with the original distribution by adaptively adjusting the temperature of the distribution.
@@ 3.3 Adaptive Watermark Temperature Scaling (AWTS)
where $\delta>0$ controls the strength of the watermark. In this way, we adaptively perturb the logit of each token by proportionally scaling up a factor of $(1+\delta\cdot\hat{v}^{(t)})$ using the obtained SLSVE. If the original logit $l_{i}^{(t)}$ is small, the $\hat{l}_{i}^{(t)}$ after the manipulation remains small, indicating that our watermark slightly perturbs these logits. Temperature scaling mainly affects tokens with large original logits $l_{i}^{(t)}$, causing the corresponding $\hat{l}_{i}^{(t)}$ to increase more significantly. In contrast to uniformly increasing the logits of tokens in the ‘green list’ as in previous work (Kirchenbauer et al., 2023a), our approach inserts a watermark by modulating the temperature of the logits. Therefore, our watermark strength adaptively depends on the original distribution, which will make the generated watermarked text more coherent. As we will show in Figure 4, AWTS also contributes to the improvement of text quality.
@@
Algorithm 2 describes the entire procedure of the proposed Adaptive Watermark algorithm. At the beginning of the text generation process, only a limited number of tokens can be used for extracting semantic embeddings, which makes the detection of the watermark unstable under attack. To maintain the stability and robustness of the watermark, we set a secret opening sentence $\hat{S}$. When the number of generated tokens is smaller than a pre-determined measure threshold $M$, we extract the semantic embedding from the opening sentence and apply SLSVE.
@@ 4 Adaptive Watermark Detection
In this section, we propose to approximate the Likelihood Ratio Test (LRT) to detect the watermark. Consider the following null hypothesis $H_{0}$ and the alternative $H_{1}$,
@@
Given a sequence of tokens $S=\{S^{(1)},S^{(2)},...,S^{(T)}\}$, we input the sequence to MM to calculate the entropy and select the potential watermarked token at each time step, following the same process as the AWTI described in Section 3.1. Denote the set of potential watermarked token as $\mathcal{W}$, i.e., for all $s^{(t)}\in\mathcal{W}$, we have $H(P_{MM}(\cdot|S_{0:t-1}))\geq\alpha$. For any $s^{(t)}\in\mathcal{W}$, if it is sampled from the perturbed distribution, the distribution would be $P_{H_{1}}(\cdot|R,S_{0:t})$. Conversely, if it is sampled from the original distribution, the distribution would be $P_{H_{0}}(\cdot|R,S_{0:t})$. Subsequently, the logits scaling vector is extracted utilizing both the Semantic Embedding Model and the Semantic Mapping Model, as detailed in Section 3.2. As we perturb the distribution with temperature scaling, the likelihood ratio can be written as
@@
where we omitted the superscript of time $t$, and in the last step, we approximated the ratio of the normalization factor in the softmax function by 1.
Figure 2: Workflow of Adaptive Detection. This figure illustrates a single step during the detection process. At each time step, we will first check if the current token is a potential watermarked token by applying AWTI to estimate the entropy of the current distribution. For potential watermarked token, the preceding text will be used to extract the logits scaling vector through SLSVE. If the corresponding value of the potential watermarked token in the logits scaling vector is positive, it will be added to the total score.
@@
Score(${\rm Score}(S=\{s^{(1)},...,s^{(T)}\})=\frac{\sum_{s^{(t)}\in\mathcal{W}}\delta\cdot\hat{v}^{(t)}_{s^{(t)}}}{|\mathcal{W}|},$ (5)
where $|\mathcal{W}|$ represents the number of all potential watermarked tokens.
5 Experiments
@@ 5.1 Experiment Setting
Implementation Details. In our watermarking algorithm, the default hyperparameters are set as follows: $\alpha=2$, $M=50$. We use Sentence-Transformers (Reimers & Gurevych, 2019) to extract the sentence embedding. The specific model we used is all-mpnet-base-v2  [^2] https://huggingface.co/sentence-transformers/all-mpnet-base-v2. More details for the Semantic Mapping Model are provided in Appendix A. We apply our method to three different models: OPT-2.7B, OPT-6.7B (Zhang et al., 2022), GPT-J-6B (Wang & Komatsuzaki, 2021), and Mistral-7B (Jiang et al., 2023) (in Appendix B.1). At each time step, the next token is sampled from the distribution using a combination of Top-$K$ and Top-$p$ sampling methods, with $K=50$ and $p=0.9$. It is worth noting that the watermarking strength $\delta$ should be chosen differently for different LLMs based on their specific capabilities. For example, we set $\delta=1.5$ for OPT-6.7B and set $\delta=0.6$ for Mistral-7B.
Baselines. Our method is compared with existing methods, including KGW-00 (Zhao et al., 2023), KGW-$1$/KGW-$2$ (Kirchenbauer et al., 2023a), EXP-edit (Kuditipudi et al., 2023). In KWG-$1$/KWG-$2$, the numeral denotes the hashing of the previous 1 or 2 tokens, respectively. For KGW-00, KGW-$1$, and KGW-$2$, the ‘green’ list token percentage is set to be $0.5$. Additional comparisons against two model non-agnostic baselines, i.e., SWEET (Lee et al., 2023) and SIR (Liu et al., 2023b), can be found in Appendix C.
Dataset and Prompt. To test the performance of our methods, the results on realnewslike subset in C4 dataset (Raffel et al., 2020) are provided here, and additional results on ELI5 (Fan et al., 2019) can be found in Appendix B.2. For C4, we use the first two sentences in each text as a prompt for LLMs and the subsequent $200$ tokens as human-generated text. For ELI5, we use the question as a prompt and the corresponding answer as human-generated text. For both datasets, we generate $200\pm 30$ tokens using the LLMs and prompts. The Semantic Mapping Model is trained using the Sentence-Compression dataset  [^3] https://huggingface.co/datasets/embedding-data/sentence-compression.
Evaluation Methods. Watermark detection performance is evaluated through the ROC-AUC value, the Best F1 Score, and the ROC curve, which illustrates the true positive rate (TPR) and false positive rate (FPR) at various thresholds. The TPR and FPR indicate the accurate identification of watermarked text and the incorrect recognition of human-written text as being watermarked. The perplexity of the text is calculated using LLama-13B (Touvron et al., 2023), measuring the quality of the text. In Appendices B.1 and B.2, we also use a more powerful LLM, GPT-3, to calculate the perplexity.
@@ 5.2 Robustness
To demonstrate the robustness of our watermark strategy, we perform a paraphrase attack on the watermarked text using two different paraphrasing methods: GPT-$3.5$ and DIPPER (Krishna et al., 2023). For GPT-$3.5$, we use the gpt-3.5-turbo-instruct version, and we employ the same prompt as used in Kirchenbauer et al. (2023b) to paraphrase the watermarked text. DIPPER is a paraphrase generation model to rewrite the text and evade the watermark detection, and the lexical diversity there is set to be $60$.
The results are summarized in Table 1, which shows that all methods demonstrate effective performance when the watermarked texts remain unaltered. However, when the watermarked text is paraphrased using GPT-$3.5$ and DIPPER, the proposed adaptive watermarking method remains high ROC-AUC and Best F1 Score, surpassing those of the baselines across different language models. The corresponding ROC curves, showcasing the performance of different watermarking methods on various language models under paraphrasing attacks, can be found in Appendix B.3. Moreover, we evaluate the TPR scores at fixed 1% and 10% FPR as detailed in Appendix B.4.
@@ 5.3 Text Quality
The quality of the generated text is evaluated by perplexity. A lower perplexity value suggests better text quality. In Figure 3, we compare the perplexity of text generated by our method with human-written text, un-watermarked text (texts generated by the model without any watermarking), and other baseline methods. Our method has a negligible impact on the perplexity compared to the un-watermarked text. It also shows comparable results with EXP-edit, which is a distortion-free watermark method. The quality of text watermarked using our approach is clearly superior to that of KGW-00 and KGW-$1$. In Appendix B.5, we also report the repetition rate of watermarked text.
@@ 5.4 Security Results
The spoofing attack (Sadasivan et al., 2023) is used to evaluate the security of our method, which is an adversarial attack designed to produce non-watermarked text that can be mistakenly identified as AI-generated by the watermark detector. This is accomplished by frequently querying the watermarked language model and identifying the potential ‘green’ list as per a particular watermarking algorithm by analyzing token frequencies in the generated text.
Specifically, for KGW-1/KGW-2, the attacker seeks to uncover the green token list for any selected fixed prefix. This entails generating 5,000 sentences under the KGW-1 and examining the frequency of tokens among 181 common tokens ($C$). From $C$, the top 50 tokens with the highest frequency ($H$) are identified. For example, the process involves analyzing the occurrence frequency of tokens following ‘the.’ As the actual green token list ($G$) is determined using the token ‘the’ as a hash, the decryption rate is then computed based on the proportion of tokens in $H$ that belong to $G$.
@@
Language Model Evaluation KGW-0 KGW-1 KGW-2 Ours
@@
For the proposed watermarking algorithm, as our ‘green’ token list (positive entries in logits scaling vector $v^{(t)}$) changes with the semantic meaning, the attacker needs to determine the list for any semantic meaning embedding. However, fixing a specific semantic is challenging because the semantics of the generated text can vary with each generation, making it difficult to measure and control consistently. Therefore, to help us qualify the decryption rate, we strengthen the spoofing attack. In our experiments, we maintain a fixed sentence embedding for all generations. This implies that the green token list remains unchanged, allowing us to verify if the green token list can be obtained for this fixed semantic embedding. We analyze the frequency of tokens in a manner similar to the previously mentioned methods.
We note that the decryption rate derived from this strengthened version of the spoofing attack serves as an upper bound because, in practice, attackers typically only have access to the watermarked LLM API, which generates watermarked text in response to input prompts. Even if an attacker repeatedly inputs the same prompt and generates multiple texts, the generated text may have different semantic meanings, and decrypting the mapping to the green token list becomes significantly more challenging.
As shown in Table 2, we can observe that KGW-0 exhibits the highest decryption rate because of its fixed ‘green/red’ list. For KGW-1 and KGW-2, the implementation of the previous token as a hash key adds a certain level of difficulty to the decryption process compared to the KGW-0. Our method demonstrates superior security performance compared to those works, even for the strengthened spoofing attack, with a decryption rate of only $0.571$. This is because the AWTI further reduces the fraction of watermarked tokens in the text, thereby diluting the evident token preference within a specific semantic range. More discussion regarding the benefits of AWTI to security can be found in Appendix D.